OpenStudy (brooke..help00):

@lowkey.s @chpatterson Solve for pi: C = 2pir. pi equals 2 times C over r pi equals C over quantity 2 times r pi equals C over r pi equals r times C over 2

OpenStudy (owlcoffee):

\[C= 2 \pi r \] When wolving for a variable, you want to isolate it, so you apply the convenient operations in order to isolate the variable, and treat the rest as normal numbers. Here, I would isolate that "r" by dividing both sides by "r". \[\frac{ C }{ r }=\frac{ 2 \pi r }{ r } \rightarrow \frac{ C }{ r }=2 \pi \]

OpenStudy (brooke..help00):

\[\pi=\frac{ C }{ R }\]